# SparseDM: Toward Sparse Efficient Diffusion Models

### v2

@@ 4.2 Result Comparison
2:4 Sparsity Results
Table 2: The comparison of 2: 4 ($50\%$ ) sparsity results. No Pruning is the method of U-ViT. U-ViT with LDM is U-ViT using Latent Diffusion Models (LDM) Rombach et al. (2022) for data process. Diff-Pruning is a structural pruning method. ASP is the method of sparse pruning from NVIDIA. DDPM Ho et al. (2020) is a diffusion model based on U-Net.
@@
U-ViT CIFAR10 32x32 No Pruning 3.20 11.34
@@
Ours 4.23 5.67
CelebA 64x64 No Pruning 2.87 11.34
@@
U-ViT with LDM MS-COCO 256x256 No Pruning 5.95 11.34
@@
Ours 8.14 5.68
MS-COCO 256x256 (Deep) No Pruning 5.48 14.72
Diff-Pruning 16.92 7.12
ASP 340.69 7.43
Ours 6.82 7.37
ImageNet 256x256 No Pruning 3.40 76.66
@@
Ours 5.43 36.84
DDPM CIFAR10 32x32 No Pruning 3.23 9.44
@@
Ours 3.66 4.87
We experimented with four datasets (CIFAR10, CelebA, MS COCO 2014, and ImageNet) and three resolutions (32$\times$32, 64$\times$64, and 256$\times$256) on Transformer-based and UNet-based DMs. We reduced the computational load by approximately $50\%$, and the FID only increased by 0.60$\sim$2.19 (1.22 on average). The higher the resolution, especially for models of similar sizes, the better the FID effect, such as U-ViT on CIFAR10 32$\times$32 and CelebA 64$\times$64. This way, our method is more suitable for model acceleration in high-resolution and high-fidelity image generation.
@@
In addition to the theoretical acceleration results of $50\%$, Table 3, we show our hardware test of the model in the basic setting. The acceleration is defined as the ratio between the running of a dense model and a sparse model in the same setting. The acceleration ratio should almost be the same across all datasets while the weights also have nothing to do with this.
In the testing code we provide, we set hyper-parameters head_dimension=1024, num_head=8, depth=1 to get the following result. We do the experiment on the two significant hyper-parameters: mlp_ratio (determines the dimension of the MLP layer according to attention layer dimension) and patch_size (determines the length of sequence for attention computation). Theoretically, only the ratio between sequence length and head dimension will impact the acceleration ratio, so we don’t bother to sweep the head dimension further.
We can see that the acceleration is among 1.1 based on the concrete parameters. It is worth mentioning that we only use the most naive acceleration of the 2:4 sparsity provided in PyTorch. Many other recent works like Hu et al. (2024) provide further acceleration, and all these kinds of methods can be integrated into our models and achieve faster training and inference.
Table 3: 2:4 ($50\%$ ) sparse model acceleration results tested on 4 A40 GPU and CIFAR10 32$\times$32. The row indices and column indices are the hyper-parameters we choose.
@@
Other Sparsity Results To better understand the performance of our sparse training method, except for the $50\%$ sparsity ratio of 2:4, we have also conducted experiments on other sparsity ratios. Since the ASP method is a sparsity tool provided by Nvidia for GPU hardware acceleration, no other sparsity ratios are provided. The following is mainly a comparison of different sparsity levels with other methods.
@@ 4.3 Ablation Study
Our method involves performing sparse mask fine-tuning on the existing trained model. Masks need to be trained during the training process. We designed three ablation experiments to demonstrate better the rationality of our method design and the necessity of each step. The first one is the STE-based method. The second one is traditional progressive sparse training of DM. The last one involves fixing the mask during the training process and not performing sparse training.
STE Sparse Training As shown in Fig. LABEL:fig:sparse_inference(d) and LABEL:fig:sparse_inference(e), on datasets CIFAR10 and CelebA64, whether it is 2:4 sparse or other sparsity ratio, the training method of STE is worse than our method, mainly because our method adds sparse mask information into the backward propagation process.
Traditional Progressive Sparse Training As shown in Fig. LABEL:fig:sparse_inference(b), DM is trained on datasets CIFAR10. Every 100,000 training steps, the sparsity of masks is increased for progressive sparse training. Training after switching sparse masks for the first time is almost ineffective. The traditional progressive sparse training does not work well on the diffusion model, especially when switching the sparsity rate every time. It is equivalent to training stopped. However, the fixed sparse mask can be continuously trained until convergence. This comparison result motivates us to propose a new method for progressively sparse training of diffusion models.
Sparse Mask Inference Unlike our trained masks, we generate untrained sparse masks for sparse pruning diffusion models. Dense model weights are imported into the sparse model, and the sparse model is not trained. Only use STE-based methods for sparse training. As shown in Fig. LABEL:fig:sparse_inference(f), on datasets CIFAR10, the result of the untrained mask is the worst because the mask did not participate in training. The STE-trained mask is worse than ours because the mask did not use progressive sparse training. At the highest sparsity rate, our method achieves an FID 137 lower than that of random masks. Our method achieves 1 less FID at the lowest sparsity rate than the untrained mask.
@@ 5.1 N: M Sparsity
A neural network with $N$:$M$ sparsity ($M$-weights contains $N$ non-zero values) satisfies that, in each group of $M$ consecutive weights of the network, there are at most $N$ weights have non-zero values Zhou et al. (2021); Chmiel et al. (2022). APEX’s Automatic SParsity (ASP) Pool and Yu (2021); Mishra et al. (2021)) is a 2:4 sparse tool provided by NVIDIA. This tool obtains a 2:4 sparse network and can achieve nearly 2x runtime speedup on NVIDIA Ampere and Hopper architecture GPUs Nvidia (2020). A provable and efficient method for finding N: M transposed masks for accelerating sparse neural training Hubara et al. (2021). Compared with dense networks, sparse network training has gradient changes, and methods such as STE Bengio et al. (2013) should be used to improve training performance. In the forward stage, sparse weight is obtained by pruning dense weight. In the backward stage, the gradient w.r.t. sparse weight will be directly applied to dense weight. SR-STE Zhang et al. (2022) enhances the backward pass information by integrating the sparse weight of the forward pass into the backward pass. DominoSearch Sun et al. (2021) found mixed N: M sparsity schemes from pre-trained dense, deep neural networks to achieve higher accuracy than the uniform-sparsity scheme with equivalent complexity constraints. The N: M sparse method is also used in SparseGPT Frantar and Alistarh (2023) to accelerate the LLM generation model GPT. Although many N: M sparse methods exist, none have been proven effective for sparse DM.
@@ 5.2 Diffusion Model Pruning
Pruning is one of the most used methods to reduce the calculation time of DNN, including DM. The pruning method was divided into structured pruning and unstructured pruning. Structural pruning aims to physically remove a group of parameters, thereby reducing the size of neural networks. In contrast, unstructured pruning involves zeroing out specific weights without altering the network structure Fang et al. (2023b). Sparsity can reduce the memory footprint of DM to fit mobile devices and shorten training time for ever-growing networks Hoefler et al. (2021). Pay attention to features and selection of useful features for the target dataset Wang et al. (2020), which can also reduce computational complexity. The parameter-grouping patterns vary widely across different models, making architecture-specific pruners, which rely on manually designed grouping schemes, non-generalizable to new architectures. Depgraph Fang et al. (2023b) can tackle general structural pruning of arbitrary architecture. Structured sparsity was also used for large language models Ma et al. (2023); Frantar and Alistarh (2023); Guo et al. (2023). Because Transformer has been proven to outperform the other networks in multiple applications, including DM. ToMe Bolya and Hoffman (2023) merge redundant tokens of Transformer. There are also proposals for early exiting for accelerated inference in DM Moon et al. (2023). Although there are many pruning methods for DM, either the sparse model FID is poor or structured 2:4 sparsity is not implemented.
@@ 6 Conclusion
In this paper, we studied how to improve the efficiency of DM by sparse matrix for 2:4 sparse acceleration GPU. The existing STE-based methods make it difficult to optimize sparse DM. To address this issue, we improve the STE method and propose a progressive approach to transfer knowledge from dense models to sparse models gradually. Our method is tested on the latest Transformer-based DM, U-ViT, and UNet-based DM, DDPM. We trained the 2:4 and other sparse models to perform better than other methods. Our approach also provides an effective solution for DM deployment on NVIDIA Ampere architecture GPUs, achieving about 1.2x acceleration.
Limitations and future work. The MACs of the sparse diffusion model are significantly reduced, but the acceleration effect on real processors is limited, especially the 2:4 sparse acceleration that can currently only reach about 1.2 times. Future work will design sparse acceleration schemes for different processors to achieve better processor acceleration.

### v3

@@ 4.2 Result Comparison
2:4 Sparsity Results.
Table 2: The comparison of 2: 4 ($50\%$ ) sparsity results. No Pruning is the method of U-ViT or DDPM. U-ViT with LDM is U-ViT using Latent Diffusion Models (LDM) (Rombach et al., 2022) for data process. Diff-Pruning is a structural pruning method. ASP is the method of sparse pruning from NVIDIA. STE-based Pruning (Bengio et al., 2013; Zhang et al., 2022) only uses sparse masks for training and inference. DDPM (Ho et al., 2020) is a diffusion model based on U-Net.
@@
U-ViT (Transformer-Based) CIFAR10 32x32 No Pruning 3.97 11.34
@@
STE-based Pruning 4.23 5.67
Ours 3.81 5.67
CelebA 64x64 No Pruning 3.30 11.34
@@
U-ViT (Transformer-Based) with LDM MS-COCO 256x256 No Pruning 5.95 11.34
@@
STE-based Pruning 8.14 5.68
Ours 7.09 5.68
ImageNet 256x256 No Pruning 3.81 76.66
@@
STE-based Pruning 5.83 36.84
Ours 5.25 36.84
DDPM (UNet-Based) CIFAR10 32x32 No Pruning 3.23 9.44
@@
STE-based Pruning 3.66 4.87
Ours 3.04 4.87
We experimented with four datasets (CIFAR10, CelebA, MS COCO 2014, and ImageNet) and three resolutions (32$\times$32, 64$\times$64, and 256$\times$256) on Transformer-based and UNet-based DMs. We reduced the computational load by approximately $50\%$, and the FID only increased by 0.44 on average. The higher the resolution, especially for models of similar sizes, the better the FID effect, such as U-ViT on CIFAR10 32$\times$32 and CelebA 64$\times$64. This way, our method is more suitable for model acceleration in high-resolution and high-fidelity image generation.
@@
Table 3: Speedup results for 2:4 ($50\%$ ) sparse model sampling on CIFAR10 32x32 data tested on 4 A40 GPUs. Patch_size and mlp_ratio are tunable hyperparameters.
@@
Based on the theoretical acceleration results of $50\%$, Table 3 shows the actual acceleration results of testing the 2:4 sparse operator on the GPU, which is approximately 1.2x. The acceleration result is defined as the ratio between the running of a dense model and a sparse model in the same setting. The acceleration ratio should almost be the same across all datasets, and the weights have nothing to do with this. There are two significant hyper-parameters of the experiment: mlp_ratio, which sets the dimension of the MLP layer according to the attention layer dimension, and patch_size, which sets the sequence length for attention computation. Based on experimental experience, the ratio of sequence length to header dimension affects the speedup ratio. Some other hyper-parameters are head_dimension is 1024, num_head is 8, and depth is 1. In the above experiments, we only used the most primitive acceleration of 2:4 sparsity provided by PyTorch. Defining more advanced acceleration operators, such as (Hu et al., 2024), provides further acceleration solutions. If we can integrate these methods into our model, our model can achieve faster training and inference.
Other Sparsity Results. To better understand the performance of our sparse training method, except for the $50\%$ sparsity ratio of 2:4, we have also conducted experiments on other sparsity ratios. Since the ASP method is a sparsity tool provided by Nvidia for GPU hardware acceleration, no other sparsity ratios are provided. The following is mainly a comparison of different sparsity levels with other methods.
@@ 4.3 Ablation Study
Our method involves performing sparse mask fine-tuning on the existing trained model. Masks need to be trained during the training process. We designed three ablation experiments to demonstrate better our method design’s rationality and the necessity of each step. The first one is the STE-based method. The second one is traditional progressive sparse training of DM. The last one involves fixing the mask during the training process and not performing sparse training.
STE Sparse Training. As shown in Fig. LABEL:fig:sparse_inference(d), on datasets CIFAR10 and CelebA64, the learning curve of STE converges significantly slower than our method, mainly because our sparse model is trained with knowledge transfer. As shown in Fig. LABEL:fig:sparse_inference(e), STE’s results are worse than ours, mainly because our method adds sparse mask information during back-propagation. At high compression ratios, such as 1:32, transfer learn sparse masks also plays an important role.
Traditional Progressive Sparse Training. As shown in Fig. LABEL:fig:sparse_inference(b), DM is trained on dataset CIFAR10. Training after switching sparse masks for the first time is almost ineffective. Every 100,000 training steps, the sparsity of masks is increased for progressive sparse training. The traditional progressive sparse training does not work well on the diffusion model, especially when switching the sparsity rate every time. It is equivalent to training stopped. However, the fixed sparse mask can be continuously trained until convergence. This comparison shows that simultaneously transforming the data and model distribution will fail DM training, prompting us to propose a new method for gradually sparsely training DMs.
Sparse Mask Inference. Unlike our trained masks, we generate untrained sparse masks for sparse pruning diffusion models. Dense model weights are imported into the sparse model, and the sparse model is not trained. As shown in Fig. LABEL:fig:sparse_inference(f), on dataset CIFAR10, the untrained mask’s result is the worst at all MACs conditions because the mask did not participate in training. The STE-trained mask is worse than ours because the mask did not transfer knowledge from dense model. Compared to the STE-trained masks, our method achieves 1 less FID at 9 sparsity rates.
@@ 5.1 N: M Sparsity
A neural network with $N$:$M$ sparsity ($M$-weights contains $N$ non-zero values) satisfies that, in each group of $M$ consecutive weights of the network, there are at most $N$ weights have non-zero values (Zhou et al., 2021; Chmiel et al., 2022). APEX’s Automatic SParsity (ASP) (Pool & Yu, 2021; Mishra et al., 2021)) is a 2:4 sparse tool provided by NVIDIA. This tool obtains a 2:4 sparse network and can achieve nearly 2x runtime speedup on NVIDIA Ampere and Hopper architecture GPUs (Nvidia, 2020). A provable and efficient method for finding N: M transposed masks for accelerating sparse neural training (Hubara et al., 2021). Compared with dense networks, sparse network training has gradient changes, and methods such as STE (Bengio et al., 2013) should be used to improve training performance. In the forward stage, sparse weight is obtained by pruning dense weight. In the backward stage, the gradient w.r.t. sparse weight will be directly applied to dense weight. SR-STE (Zhang et al., 2022) enhances the backward pass information by integrating the sparse weight of the forward pass into the backward pass. DominoSearch (Sun et al., 2021) found mixed N: M sparsity schemes from pre-trained dense, deep neural networks to achieve higher accuracy than the uniform-sparsity scheme with equivalent complexity constraints. The N: M sparse method is also used in SparseGPT (Frantar & Alistarh, 2023) to accelerate the LLM generation model GPT. Although many N: M sparse methods exist, none have been proven effective for sparse DM.
@@ 5.2 Diffusion Model Pruning
Pruning is one of the most used methods to reduce the calculation time of DNN, including DM. The pruning method was divided into structured pruning and unstructured pruning. Structural pruning aims to physically remove a group of parameters, thereby reducing the size of neural networks. In contrast, unstructured pruning involves zeroing out specific weights without altering the network structure (Fang et al., 2023a). Sparsity can reduce the memory footprint of DM to fit mobile devices and shorten training time for ever-growing networks (Hoefler et al., 2021). Pay attention to features and selection of useful features for the target dataset (Wang et al., 2020), which can also reduce computational complexity. The parameter-grouping patterns vary widely across different models, making architecture-specific pruners, which rely on manually designed grouping schemes, non-generalizable to new architectures. Depgraph (Fang et al., 2023a) can tackle general structural pruning of arbitrary architecture. Structured sparsity was also used for large language models (Ma et al., 2023; Frantar & Alistarh, 2023; Guo et al., 2023). Because Transformer has been proven to outperform the other networks in multiple applications, including DM. ToMe (Bolya & Hoffman, 2023) merge redundant tokens of Transformer. There are also proposals for early exiting for accelerated inference in DM (Moon et al., 2023). Although there are many pruning methods for DM, either the sparse model FID is poor or structured 2:4 sparsity is not implemented.
@@ 6 Conclusion
In this paper, we studied how to improve the efficiency of DM by sparse matrix for 2:4 sparse acceleration GPU. The existing STE-based methods make it difficult to optimize sparse DM. To address this issue, we improve the STE method and propose to gradually transfer knowledge from dense models to sparse models. Our method is tested on the latest Transformer-based DM, U-ViT, and UNet-based DM, DDPM. We trained the 2:4 and other sparse models to perform better than other methods. Our approach also provides an effective solution for DM deployment on NVIDIA Ampere architecture GPUs, achieving about 1.2x acceleration.
7 Ethics Statement
While our motivation is to make diffusion models generate images faster, we also recognize the ethical implications inherent in model pruning, as sparse optimization may introduce unintended biases. There are differences between traditional evaluation metrics (such as FID scores) and human eye evaluation, and optimization for FID may not capture subtle deviations in human vision. In addition to performance improvements to diffusion models, we are also actively focusing on the ethical aspects of AI deployment. We are committed to preventing the spread of content that may have a negative social impact. Prevent unintended consequences by acknowledging the limitations of FID indicators.
8 Reproducibility
The anonymous address of the source code for the paper is https://github.com/aaa-bbb-111/SparseDM. The paper code implementation consists of two parts. The first part, sparse training, trains the sparse DM and verifies the acceleration theory of the sparse diffusion model. The second part, GPU inference, verifies GPU inference acceleration.
Appendix A appendix
Limitations and Future Work. The MACs of the sparse diffusion model are significantly reduced, but the acceleration effect on real processors is limited, especially the 2:4 sparse acceleration that can currently only reach about 1.2 times. Future work will design sparse acceleration schemes for different processors to achieve better acceleration.